OpenStudy (anonymous):

Can someone help please!!!! ill give medals!!!! Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. Function Transformation from f(x) f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@phi

OpenStudy (phi):

to graph a function you need to know the details. You did not post them.

OpenStudy (anonymous):

sorry im really confused, this is the function im suppose to be graphing f(x) = 27(1.3)x but i dont understand how im suppose to graph them like this f(x + 2) f(–x) f(x) + 3

OpenStudy (phi):

I use geogebra to plot things.

OpenStudy (phi):

you can download geogebra and try it. or you can find (x,y) pairs (use a calculator) and plot them on graph paper

OpenStudy (anonymous):

oh ok, thxs but im still confused on how to write the explain the effect of the transformations.

OpenStudy (phi):

f(x + 2) means replace x with (x+2) in the "definition" \[ f(x) = 27(1.3)^x \] becomes \[ g(x) = 27(1.3)^{x+2} \] (we rename the function because it is different) Here they are

OpenStudy (phi):

f(–x) means we replace x with -x \[ h(x) = 27(1.3)^{-x} \]

OpenStudy (phi):

f(x) + 3 that means add 3 to every value (it makes the new function 3 higher than the original)

OpenStudy (anonymous):

oh ok!! thank you so much that helped alot!!